OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. A(5, 5), m = 3

OpenStudy (amistre64):

standard form is described by: ax+by = c

OpenStudy (amistre64):

where slope = -a/b

OpenStudy (amistre64):

another method is just to put it in slope intercept form and manipultate it into standard form

OpenStudy (amistre64):

y = mx -mPx+Py y = 3x -3(5) + 5 y = 3x -15+5 y = 3x -10 ; now change it to standard form :)